# Directional Convergence, Benign Overfitting of Gradient Descent in leaky ReLU two-layer Neural Networks

### v1

@@ Appendix A Proof of Directional Convergence
×$\qquad\times\Big{\{}\sqrt{q_{-}+\gamma^{2}q_{+}}\big{\{}(q_{+}+\gamma^{2}q_{-}-\gamma)n_{+}\|\boldsymbol{\mu}\|^{2}+(q_{+}+\gamma^{2}q_{-})R\big{\}}$
@@
Under regime (ii), for $i\in I_{+}$ we have
$d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})$ $\leq q_{\gamma}(n_{+}+\|\boldsymbol{\mu}\|^{2})(n_{-}+\|\boldsymbol{\mu}\|^{2})$ (48)
$\leq q_{\gamma}\left(\frac{1+\tilde{\varepsilon}_{3}}{2}n\|\boldsymbol{\mu}\|^{2}+R\right)^{2}$
$\leq q_{\gamma}\left(\frac{(1+\tilde{\varepsilon}_{3})\lambda}{2}+1\right)^{2}R^{2}.$
We also have
$(q_{-}+\gamma^{2}q_{+}-\gamma)n_{-}\|\boldsymbol{\mu}\|^{2}+(q_{-}+\gamma^{2}q_{+})R$ (49)
$\geq$ $q_{\gamma}R-\frac{\gamma(1-\gamma)(1+\tilde{\varepsilon}_{3})}{2}n\|\boldsymbol{\mu}\|^{2}$
$>$ $\frac{q_{\gamma}}{2}R.$
By (46), (48), and (49), we have
$\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}y_{i}\boldsymbol{e}_{i}>B_{ii}^{-1}R^{-1}\left[\frac{q_{\gamma}^{\frac{1}{2}}}{2\big{\{}\frac{(1+\tilde{\varepsilon}_{3})\lambda}{2}+1\big{\}}^{2}}-2q_{\gamma}^{-3/2}\sqrt{n}\tilde{\varepsilon}\right]\geq 0.$ (50)
The same conclusion analogously holds for $i\in I_{-}$ as well.
Under regime (iii), we have that
$d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})$ (51)
$=$ $(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\{n_{+}\|\boldsymbol{\mu}\|^{2}+R\}\{n_{-}\|\boldsymbol{\mu}\|^{2}+R\}-\gamma^{2}n_{+}n_{-}\|\boldsymbol{\mu}\|^{4}$
$=$ $(1-\gamma^{2})^{2}q_{+}q_{-}n_{+}n_{-}\|\boldsymbol{\mu}\|^{4}+(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\{n\|\boldsymbol{\mu}\|^{2}+R\}R$
$\leq$ $\Big{[}\frac{(1-\gamma^{2})^{2}}{16}n^{2}\|\boldsymbol{\mu}\|^{4}+q_{\gamma}\{n\|\boldsymbol{\mu}\|^{2}+R\}R\Big{]}$
$\leq$ $\Big{[}\frac{(1-\gamma^{2})^{2}}{16}+q_{\gamma}(1+\lambda^{-1})\lambda^{-1}\Big{]}(n\|\boldsymbol{\mu}\|^{2})^{2}.$
With this and (46), we have for $i\in I_{+}$
@@
$>$ $B_{ii}^{-1}\left[\frac{q_{\gamma}^{\frac{1}{2}}\{(q_{\gamma}-\gamma)(1-\tilde{\varepsilon}_{3})\frac{n}{2}\|\boldsymbol{\mu}\|^{2}\}}{\Big{\{}\frac{(1-\gamma^{2})^{2}}{16}+q_{\gamma}(1+\lambda^{-1})\lambda^{-1}\Big{\}}(n\|\boldsymbol{\mu}\|^{2})^{2}}-2q_{\gamma}^{-\frac{3}{2}}\sqrt{n}\tilde{\varepsilon}R^{-1}\right]$
$=$ $(B_{ii}n\|\boldsymbol{\mu}\|^{2})^{-1}q_{\gamma}^{-\frac{3}{2}}\left[\frac{q_{\gamma}^{2}(q_{\gamma}-\gamma)(1-\tilde{\varepsilon}_{3})}{2\Big{\{}\frac{(1-\gamma^{2})^{2}}{16}+q_{\gamma}(1+\lambda^{-1})\lambda^{-1}\Big{\}}}-2\sqrt{n}\tilde{\varepsilon}\frac{n\|\boldsymbol{\mu}\|^{2}}{R}\right]\geq 0.$
@@
[…]
@@
By Lemma A.2, the network $f(x;\hat{W})$ has the linear decision boundary defined by
@@
$\boldsymbol{\bar{w}}$ $=\left(q_{+}X^{\top}B_{+}+q_{-}X^{\top}B_{-}\right)(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}$
@@
$\|\boldsymbol{\bar{w}}\|^{2}$ $=\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}\left(\sum_{j\in\{+,-\}}q_{j}B_{j}X\right)\left(\sum_{j\in\{+,-\}}q_{j}B_{j}X\right)^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y},$ (55)
$\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle$ $=\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}\left(q_{+}B_{+}X+q_{-}B_{-}X\right)\boldsymbol{\mu}$
@@
We need to upper bound $\|\boldsymbol{\bar{w}}\|^{2}$ and lower bound $\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle$, respectively, in order to later establish a classification error bound of $\boldsymbol{\bar{w}}$.
To obtain an upper bound for $\|\boldsymbol{\bar{w}}\|^{2}$, the following inequality is useful.
Lemma B.7.
$\|\boldsymbol{\bar{w}}\|^{2}\leq 2\max\{q_{+},q_{-}\}\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}.$
@@
Since $(U+V)(U+V)^{\top}\succeq O$ and $(U-V)(U-V)^{\top}\succeq O$, we have
$2(UU^{\top}+VV^{\top})=(U+V)(U+V)^{\top}+(U-V)(U-V)^{\top}\succeq(U+V)(U+V)^{\top}.$
Letting $U=q_{+}B_{+}X$ and $V=q_{-}B_{-}X$, we have by (55)
$\|\boldsymbol{\bar{w}}\|^{2}$ $\leq 2\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}(q_{+}^{2}B_{+}XX^{\top}B_{+}+q_{-}^{2}B_{-}XX^{\top}B_{-})(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}$
$\leq 2\max\{q_{+},q_{-}\}\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}(q_{+}B_{+}XX^{\top}B_{+}+q_{-}B_{-}XX^{\top}B_{-})(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}$
$=2\max\{q_{+},q_{-}\}\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}.$
Lemma B.8.
Suppose all the assumptions of Lemma B.6 hold with $\tilde{\varepsilon}_{3}\leq\tfrac{1}{2}$ and $\lambda\geq\frac{\gamma}{4}$ in (iii). Then, there is a constant $\tilde{c}_{1}$ which only depends on $\gamma$ and $q_{+}$ such that we have
@@
By Lemma B.7, it is reduced to obtaining an upper bound of $\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}$. By Lemma B.5 and the definition of $q_{\gamma}$, we have
$\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}$ $\leq(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}\boldsymbol{y}+\frac{2q_{\gamma}^{-1}\tilde{\varepsilon}}{R}\left(\frac{n_{+}}{q_{+}+\gamma^{2}q_{-}}+\frac{n_{-}}{q_{-}+\gamma^{2}q_{+}}\right)$ (57)
$\leq(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}\boldsymbol{y}+\frac{2q_{\gamma}^{-2}n\tilde{\varepsilon}}{R}.$
@@
Case $q_{+}=1,q_{-}=0$: By (36), we have
$(B_{+}^{-1}\boldsymbol{y})^{\top}A^{-1}(B_{+}^{-1}\boldsymbol{y})$ (58)
@@
$\qquad\qquad\left.+\gamma^{-1}\left(\gamma^{-1}-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{n_{+}+\gamma^{-1}n_{-}}{n}\right)n_{-}\right\}.$
By (57), we have under the regime (i)
$\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}\;\leq$ $\;R^{-1}\left\{\left(1-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{n_{+}+\gamma^{-1}n_{-}}{n}\right)n_{+}\right.$ (59)
@@
$\leq R^{-1}\left\{(n_{+}+\gamma^{-2}n_{-})+2nq_{\gamma}^{-2}\tilde{\varepsilon}\right\}$
$\leq\left(\gamma^{-2}+2q_{\gamma}^{-2}\tilde{\varepsilon}\right)\frac{n}{R}$
$\leq\frac{3}{2\gamma^{2}}\frac{n}{R},$
where the last inequality is due to the condition on $\tilde{\varepsilon}$ of the regime (i).
Case $q_{+}=0,q_{-}=1$: Similarly, by (39) and (57), we obtain the same bound as the case $q_{+}=1,q_{-}=0$.

### v2

@@ Appendix B Proof of Directional Convergence
$\qquad\times\Big\{\sqrt{q_{-}+\gamma^{2}q_{+}}\big\{(q_{+}+\gamma^{2}q_{-}-\gamma)n_{+}\|\boldsymbol{\mu}\|^{2}+(q_{+}+\gamma^{2}q_{-})R\big\}$
@@
Under regime (ii), (43) implies
$d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\leq\frac{5q_{\gamma}}{4}(1+\lambda)^{2}R^{2}.$ (53)
By $q_{\gamma}>\gamma^{2}$, we have
$(q_{-}+\gamma^{2}q_{+}-\gamma)n_{-}\|\boldsymbol{\mu}\|^{2}+(q_{-}+\gamma^{2}q_{+})R$ $>q_{\gamma}R-\gamma(1-\gamma)n_{-}\|\boldsymbol{\mu}\|^{2}$ (54)
$\geq q_{\gamma}R-\gamma(1-\gamma)n\|\boldsymbol{\mu}\|^{2}$
$\geq\frac{q_{\gamma}}{2}R,$
where the last inequality is due to the assumption on $\lambda$.
Similarly, we have
$(q_{+}+\gamma^{2}q_{-}-\gamma)n_{+}\|\boldsymbol{\mu}\|^{2}+(q_{+}+\gamma^{2}q_{-})R>\frac{q_{\gamma}}{2}R.$ (55)
Therefore, for $i\in I_{+}$, (51), (53), and (54) imply
$\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}y_{i}\boldsymbol{e}_{i}>B_{ii}^{-1}R^{-1}\left[\frac{2q_{\gamma}^{\frac{1}{2}}}{5(1+\lambda)^{2}}-2q_{\gamma}^{-\frac{3}{2}}\sqrt{n}\tilde{\varepsilon}\right]\geq 0,$ (56)
where the last inequality is due to the assumption on $\sqrt{n}\tilde{\varepsilon}$.
The same conclusion analogously holds for $i\in I_{-}$ as well by (52), (53), and (54).
Under regime (iii), by (43), we have that
$d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})$ $\leq\frac{5q_{\gamma}}{4}(1+\lambda^{-1})^{2}\left(n\|\boldsymbol{\mu}\|^{2}\right)^{2}.$ (57)
With (51) and (57), we have for $i\in I_{+}$
@@
$>$ $B_{ii}^{-1}\left[\frac{q_{\gamma}^{\frac{1}{2}}\{(q_{\gamma}-\gamma)(1-\tilde{\varepsilon}_{3})\frac{n}{2}\|\boldsymbol{\mu}\|^{2}\}}{\frac{5q_{\gamma}}{4}(1+\lambda^{-1})^{2}(n\|\boldsymbol{\mu}\|^{2})^{2}}-2q_{\gamma}^{-\frac{3}{2}}\sqrt{n}\tilde{\varepsilon}R^{-1}\right]$
$=$ $(B_{ii}n\|\boldsymbol{\mu}\|^{2})^{-1}q_{\gamma}^{-\frac{3}{2}}\left[\frac{q_{\gamma}^{2}(q_{\gamma}-\gamma)(1-\tilde{\varepsilon}_{3})}{5(1+\lambda^{-1})^{2}}-2\sqrt{n}\tilde{\varepsilon}\frac{n\|\boldsymbol{\mu}\|^{2}}{R}\right]\geq 0.$
@@
By similar argument we can show under (iii), an upper bound of the similar form as (58) for $i\in I_{+}$ whose dominant term is $(q_{-}+\gamma^{2}q_{+}-\gamma)n\|\boldsymbol{\mu}\|^{2}$. Similarly, for $i\in I_{-}$, we have an upper bound whose dominant term is $(q_{+}+\gamma^{2}q_{-}-\gamma)n\|\boldsymbol{\mu}\|^{2}$. If we have $q_{\gamma}<\gamma$, exactly one of $q_{-}+\gamma^{2}q_{+}-\gamma$ and $q_{+}+\gamma^{2}q_{-}-\gamma$ becomes negative and the other stays positive. Suppose $q_{-}+\gamma^{2}q_{+}-\gamma<0$, then we have $\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}y_{i}\boldsymbol{e}_{i}<0$ for all $i\in I_{+}$ when $n\|\boldsymbol{\mu}\|^{2}$ is sufficiently larger than $R$, which means the equivalence of $\boldsymbol{\hat{w}}=\tilde{X}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}$ provably fails in such case. The same applies with the case $q_{+}+\gamma^{2}q_{-}-\gamma<0$.
@@
[…]
@@
$\boldsymbol{\bar{w}}$ $=X^{\top}\left(q_{+}B_{+}+q_{-}B_{-}\right)(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}.$
@@
$\|\boldsymbol{\bar{w}}\|^{2}$ $=\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}\left(q_{+}B_{+}+q_{-}B_{-}\right)XX^{\top}\left(q_{+}B_{+}+q_{-}B_{-}\right)(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y},$ (61)
$\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle$ $=\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}\left(q_{+}B_{+}+q_{-}B_{-}\right)X\boldsymbol{\mu}$
@@
We need to obtain bounds of $\|\boldsymbol{\bar{w}}\|^{2}$ and $\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle$, respectively, in order to later establish a classification error bound of $\boldsymbol{\bar{w}}$.
To obtain bounds of $\|\boldsymbol{\bar{w}}\|^{2}$, the following inequality is useful.
Lemma C.9.
Suppose that the conclusions of Lemma C.2, C.5, and C.8 hold. Then, we have
$\left|\|\boldsymbol{\bar{w}}\|^{2}-(\|\boldsymbol{\mu}\|^{2}K_{1}+RK_{2})\right|\leq\tilde{\varepsilon}RK_{2},$
and
$\left|\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle-\|\boldsymbol{\mu}\|^{2}\sqrt{K_{1}}\right|\leq\frac{q_{\gamma}^{-\frac{1}{2}}}{3}\tilde{\varepsilon}_{2}\|(B^{-1}y)^{\top}A^{-1}\|R+\frac{2q_{\gamma}^{-2}}{3}\tilde{\varepsilon}\cdot\sqrt{n}\tilde{\varepsilon}_{2},$
where
$K_{1}$ $:=\left\{\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}(q_{+}B_{+}+q_{-}B_{-})\boldsymbol{y}\right\}^{2}$
$=\left\{(B^{-1}\boldsymbol{y})^{\top}B(\tilde{X}\tilde{X}^{\top})^{-1}B\cdot B^{-1}(q_{+}B_{+}+q_{-}B_{-})\boldsymbol{y}\right\}^{2}\text{ and}$ (63)
$K_{2}$ $:=\left\|(q_{+}B_{+}+q_{-}B_{-})(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}\right\|^{2}$
$=\left\|(q_{+}B_{+}+q_{-}B_{-})B^{-1}\cdot B(\tilde{X}\tilde{X}^{\top})^{-1}B\cdot B^{-1}\boldsymbol{y}\right\|^{2}.$ (64)
@@
Letting $\boldsymbol{v}=\left(q_{+}B_{+}+q_{-}B_{-}\right)(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}$, we have $\|\boldsymbol{\bar{w}}\|^{2}=\boldsymbol{v}^{\top}XX^{\top}\boldsymbol{v}$.
Then, by (61) and the conclusion of Lemma C.2, we have
$\left|\|\boldsymbol{\bar{w}}\|^{2}-\left(\|\boldsymbol{\mu}\|^{2}\langle\boldsymbol{v},\boldsymbol{y}\rangle^{2}+R\|\boldsymbol{v}\|^{2}\right)\right|\leq\tilde{\varepsilon}R\|\boldsymbol{v}\|^{2}.$
The first desired inequality now follows by noting $\langle\boldsymbol{v},\boldsymbol{y}\rangle^{2}=K_{1}$ and $\|\boldsymbol{v}\|^{2}=K_{2}$.
To show the second desired inequality, we note that (62) implies
$\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle$ $=\|\boldsymbol{\mu}\|^{2}\sqrt{K_{1}}+(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}(q_{+}B_{+}+q_{-}B_{-})Z\boldsymbol{\mu}$ (65)
$\qquad\qquad\qquad+(B^{-1}\boldsymbol{y})^{\top}\left[B(\tilde{X}\tilde{X}^{\top})^{-1}-A^{-1}\right]B^{-1}(q_{+}B_{+}+q_{-}B_{-})Z\boldsymbol{\mu}.$
By Lemma C.5 and (65), we have
$\left|\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle-\|\boldsymbol{\mu}\|^{2}\sqrt{K_{1}}\right|$ $\leq\|(B^{-1}\boldsymbol{y})^{\top}A^{-1}\|\|B^{-1}\|\|q_{+}B_{+}+q_{-}B_{-}\|\|Z\boldsymbol{\mu}\|$
$\qquad\qquad+\|B^{-1}\|\|\boldsymbol{y}\|\frac{2q_{\gamma}^{-1}\tilde{\varepsilon}}{R}\|B^{-1}\|\|q_{+}B_{+}+q_{-}B_{-}\|\|\|Z\boldsymbol{\mu}\|$
$\leq\frac{q_{\gamma}^{-\frac{1}{2}}}{3}\tilde{\varepsilon}_{2}\|(B^{-1}\boldsymbol{y})^{\top}A^{-1}\|R+\frac{2q_{\gamma}^{-2}}{3}\tilde{\varepsilon}\cdot\sqrt{n}\tilde{\varepsilon}_{2},$
where the second inequality follows from the definition of $\tilde{\varepsilon}_{2}$, $\|B^{-1}\|\leq q_{\gamma}^{-\frac{1}{2}}$, and $\|q_{+}B_{+}+q_{-}B_{-}\|\leq 1$.
Lemma C.10.
Suppose that all the assumptions of Lemma C.8 hold with $\tilde{\varepsilon}_{3}\leq\frac{1}{2}$ in (ii) and (iii). Then there exists constant $\tilde{c}_{1}$ which depends only on $\gamma$ and $q_{+}$ such that
@@
Furthermore, there exist constants $\hat{c}_{1}$ and $N$ which only depends on $\gamma$, $q_{+}$ (and $\lambda$ under (ii) and (iii)) such that $n\geq N$ implies
$\|\boldsymbol{\bar{w}}\|^{2}\geq\hat{c}_{1}\frac{n}{n\|\boldsymbol{\mu}\|^{2}+R}.$
@@
By Lemma C.9, the proof is reduced to bounding $K_{i},i=1,2$.
By Lemma C.5, we have
$\left|(B^{-1}\boldsymbol{y})^{\top}B(\tilde{X}\tilde{X}^{\top})^{-1}B\cdot B^{-1}(q_{+}B_{+}+q_{-}B_{-})\boldsymbol{y}-(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}(q_{+}B_{+}+q_{-}B_{-})\boldsymbol{y}\right|$ (66)
$\leq$ $\frac{2q_{\gamma}^{-1}\tilde{\varepsilon}}{R}\|B^{-1}\boldsymbol{y}\|\|B^{-1}(q_{+}B_{+}+q_{-}B_{-})\boldsymbol{y}\|$
$\leq$ $\frac{2q_{\gamma}^{-1}\tilde{\varepsilon}}{R}\left\|\frac{1}{\sqrt{q_{+}+\gamma^{2}q_{-}}}\boldsymbol{1}_{+}-\frac{1}{\sqrt{q_{-}+\gamma^{2}q_{+}}}\boldsymbol{1}_{-}\right\|\cdot\left\|\frac{q_{+}+\gamma q_{-}}{\sqrt{q_{+}+\gamma^{2}q_{-}}}\boldsymbol{1}_{+}-\frac{q_{-}+\gamma q_{+}}{\sqrt{q_{-}+\gamma^{2}q_{+}}}\boldsymbol{1}_{-}\right\|$
$\leq$ $\frac{2q_{\gamma}^{-1}\tilde{\varepsilon}}{R}\sqrt{\frac{1}{q_{+}+\gamma^{2}q_{-}}n_{+}+\frac{1}{q_{-}+\gamma^{2}q_{+}}n_{-}}\cdot\sqrt{\frac{(q_{+}+\gamma q_{-})^{2}}{q_{+}+\gamma^{2}q_{-}}n_{+}+\frac{(q_{-}+\gamma q_{+})^{2}}{q_{-}+\gamma^{2}q_{+}}n_{-}}$
$\leq$ $\frac{2q_{\gamma}^{-1}\tilde{\varepsilon}}{R}\sqrt{\frac{n}{q_{\gamma}}}\cdot\sqrt{\frac{n}{q_{\gamma}}}$
$\leq$ $\frac{2q_{\gamma}^{-2}n\tilde{\varepsilon}}{R}.$
Similarly, we have
[…]
$\leq$ $\frac{2q_{\gamma}^{-1}\tilde{\varepsilon}}{R}\|(q_{+}B_{+}+q_{-}B_{-})B^{-1}\|\cdot\|B^{-1}\boldsymbol{y}\|$
$\leq$ $\frac{2q_{\gamma}^{-1}\tilde{\varepsilon}}{R}\max\left\{\frac{q_{+}+\gamma q_{-}}{\sqrt{q_{+}+\gamma^{2}q_{-}}},\frac{q_{-}+\gamma q_{+}}{\sqrt{q_{-}+\gamma^{2}q_{+}}}\right\}\cdot\sqrt{\frac{n}{q_{\gamma}}}$
$\leq$ $\frac{2q_{\gamma}^{-1}\tilde{\varepsilon}}{R}\frac{1}{\sqrt{q_{\gamma}}}\cdot\sqrt{\frac{n}{q_{\gamma}}}$
$\leq$ $\frac{2q_{\gamma}^{-2}\sqrt{n}\tilde{\varepsilon}}{R}.$
@@
Regime (i): We first note that $n\|\boldsymbol{\mu}\|^{2}<\frac{\gamma}{1+(1+\gamma)\tilde{\varepsilon}_{3}}R\leq\gamma R$ implies
$R^{-1}\leq(1+\gamma)(n\|\boldsymbol{\mu}\|^{2}+R)^{-1}.$ (68)
We first consider the case when $q_{+}=1,q_{-}=0$. By (33), we have
$(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}(q_{+}B_{+}+q_{-}B_{-})\boldsymbol{y}$ (69)
$=$ $(B_{+}^{-1}\boldsymbol{y})^{\top}A^{-1}(B_{+}^{-1}\boldsymbol{y})$
@@
$\qquad\qquad\left.+\gamma^{-1}\left(\gamma^{-1}-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{n_{+}+\gamma^{-1}n_{-}}{n}\right)n_{-}\right\}$
$\leq$ $\frac{n_{+}+\gamma^{-2}n_{-}}{R}$
$\leq$ $(1+\gamma)\gamma^{-2}\frac{n}{n\|\boldsymbol{\mu}\|^{2}+R},$
where the last line is due to (68).
We also have
$(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}(q_{+}B_{+}+q_{-}B_{-})\boldsymbol{y}$ (70)
$\geq$ $R^{-1}\left\{\left(1-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{(1+\tilde{\varepsilon}_{3})(1+\gamma^{-1})}{2}\right)n_{+}\right$.
$\qquad\qquad\left.+\gamma^{-1}\left(\gamma^{-1}-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{(1+\tilde{\varepsilon}_{3})(1+\gamma^{-1})}{2}\right)n_{-}\right\}$
$\geq$ $R^{-1}\left\{\left(1-\frac{1}{2}\right)n_{+}+\gamma^{-1}\left(\gamma^{-1}-\frac{1}{2}\right)n_{-}\right\}$
$\geq$ $\{2(1+\gamma)\}^{-1}\frac{n}{n\|\boldsymbol{\mu}\|^{2}+R},$
where the first inequality is due to $\tilde{E}_{3}$, the second inequality is due to the assumption $n\|\boldsymbol{\mu}\|^{2}<\frac{\gamma}{1+(1+\gamma)\tilde{\varepsilon}_{3}}R$, and the last inequality is due to $\gamma\in(0,1)$.
Analogously, we have the same results for $q_{+}=0,q_{-}=1$.
With these, (66) with (68), and noting that $q_{\gamma}=\gamma^{2}$ when $q_{+}q_{-}=0$, we have under the regime (i)
$\left\{\frac{3}{8(1+\gamma)}\right\}^{2}\frac{n^{2}}{(n\|\boldsymbol{\mu}\|^{2}+R)^{2}}$ $\leq\left(\frac{1}{2(1+\gamma)}-2(1+\gamma)\gamma^{-4}\tilde{\varepsilon}\right)^{2}\frac{n^{2}}{(n\|\boldsymbol{\mu}\|^{2}+R)^{2}}$ (71)
$\leq K_{1}\leq(1+\gamma)^{2}\left(\gamma^{-2}+2\gamma^{-4}\tilde{\varepsilon}\right)^{2}\frac{n^{2}}{(n\|\boldsymbol{\mu}\|^{2}+R)^{2}}$
for sufficiently large $n$.
Similarly, (33) implies
@@
[…]
$=$ $R^{-1}\left\{\left(1-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{n_{+}+\gamma^{-1}n_{-}}{n}\right)^{2}n_{+}\right$.
$\qquad\qquad\left.+\gamma^{-2}\left(\gamma^{-1}-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{n_{+}+\gamma^{-1}n_{-}}{n}\right)^{2}n_{-}\right\}^{\frac{1}{2}},$
and so under the regime (i), we have
$\frac{1}{2(1+\gamma)}\frac{\sqrt{n}}{n\|\boldsymbol{\mu}\|^{2}+R}\leq\left\|(q_{+}B_{+}+q_{-}B_{-})B^{-1}\cdot A^{-1}\cdot B^{-1}\boldsymbol{y}\|\right|\leq(1+\gamma)\gamma^{-2}\frac{\sqrt{n}}{n\|\boldsymbol{\mu}\|^{2}+R}.$ (73)
With (67), we have
$\left\{\frac{3}{8(1+\gamma)}\right\}^{2}\frac{n}{(n\|\boldsymbol{\mu}\|^{2}+R)^{2}}$ $\leq\left(\frac{1}{2(1+\gamma)}-2(1+\gamma)\gamma^{-4}\tilde{\varepsilon}\right)^{2}\frac{n}{R^{2}}$ (74)
$\leq K_{2}\leq(1+\gamma)^{2}\left(\gamma^{-2}+2\gamma^{-4}\tilde{\varepsilon}\right)^{2}\frac{n}{(n\|\boldsymbol{\mu}\|^{2}+R)^{2}}$
for sufficiently large $n$.
Thus, by Lemma C.9, (71), and (74), we have
$\left[\left\{\frac{3}{8(1+\gamma)}\right\}^{2}-(1+\gamma)^{2}\left(\gamma^{-2}+2\gamma^{-4}\tilde{\varepsilon}\right)^{2}\tilde{\varepsilon}\right]\frac{n}{n\|\boldsymbol{\mu}\|^{2}+R}$
$\qquad\qquad\qquad\leq\|\boldsymbol{\bar{w}}\|^{2}\leq(1+\tilde{\varepsilon})(1+\gamma)^{2}\left(\gamma^{-2}+2\gamma^{-4}\tilde{\varepsilon}\right)^{2}\frac{n}{n\|\boldsymbol{\mu}\|^{2}+R},$
with the left hand side being positive for sufficiently large $n$.